OpenStudy (anonymous):

a horizontal force of 600.N.is used to push a box 8 m across a room. which variables must b known to determine the power used in moving the box.

OpenStudy (anonymous):

\[Power = \frac{ Work }{ time}=\frac{ Force \times distance }{ time}\]You already have two of the variables what's the third?

OpenStudy (anonymous):

dpower (j/s)=dw(j)/ds(t)---> intgration (p)dp=T intigration(dw) and know that: W(j)=F(n).D(m)cos teta. P=\[\frac{ w }{ t}\]